OpenStudy (anonymous):

PERCENT OF CHANGE? 12,200 to 13,908

OpenStudy (anonymous):

1.14 = 114%

jimthompson5910 (jim_thompson5910):

Percent Change= [(New - Old)/Old] * 100 Percent Change= [(13908-12200)/12200] * 100 Percent Change= [1708/12200] * 100 Percent Change= 0.14 * 100 Percent Change= 14% Since the percent change is positive, this is a 14% increase

OpenStudy (anonymous):

13,908 - 12,200 = 1708 1708 / 12,200 = 0.14 0.14 * 100 = 14%

OpenStudy (anonymous):

14 or 114

jimthompson5910 (jim_thompson5910):

14%